OpenStudy (anonymous):

What is the scientific notation for the number of kilometers light will travel in 60 seconds? (time 1sec) (distance 186,000mi 300,000 km)

OpenStudy (anonymous):

1.8 x 10^7 km/minute

OpenStudy (anonymous):

anserws 5.0 * 10(to the forth power) 3.0* 10(to the sixth power) 1.116*10(to the senth power) or 1.8*10 ( to the seventh power)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

When doing scientific notation you need to count the number of place you move the decimal point. If you move it to the left it's positive. If you move it to the right it's negative. The number o fplaces moved will be the 10^x. So in this case we move 5 so we have: \[186000*60=1.116*10^7miles\] \[300000*60=1.8*10^7km\]

OpenStudy (anonymous):

ok cool, thanks

OpenStudy (anonymous):

I meant moved 7 not 5. Hurray for typos ;)